# Quantum Gate Decomposition

### v1

@@ Abstract.
Quantum Computing, Quantum Programming, Quantum Compiler, Gade Decomposition, Ket
copyright: none
@@ 1. Introduction
For programming languages and platforms that treat the quantum bit (qubit) as a first-class citizen in quantum programming (Svore et al., 2018; Da Rosa and De Santiago, 2022), coding a quantum application involves invoking functions known as quantum gates. Quantum gates have no side effects on the classical state of the program; they only affect the quantum state and can induce superposition and entanglement (Nielsen and Chuang, 2010). Another class of functions, known as measurements, is used to extract information from the quantum state, returning it as classical data while causing the collapse of the quantum state, i.e., destroying the superposition.
Figure 1(a) illustrates a simple quantum program written in Python using the Ket quantum programming platform (Da Rosa and De Santiago, 2022). This code implements the Grover diffusion operator, a key component of Grover’s quantum search algorithm (Grover, 1996), utilizing the Hadamard (H), Pauli-X (X), and Pauli-Z (Z) gates. The term quantum gate, or more precisely quantum logical gate, is inspired by classical circuit and logic gates. Each quantum code has an equivalent quantum circuit; for example, the circuit in Figure 1(b) is equivalent to the code in Figure 1(a) applied for 4-qubits.
@@
{minted}
[frame=lines,fontsize=,breaklines]py def diffusion(q : Quant): with around(cat(H, X), q): with control(q[:-1]): Z(q[-1])
@@
Figure 1. Grover diffusion operation implemented using Ket (LABEL:sub@fig:cirq:diffusion), along with its corresponding high-level (LABEL:sub@fig:cirq:diffusion) and decomposed (LABEL:sub@fig:cirq:diffusion_decompose) quantum circuit for 4 qubits.
@@
The remainder of this paper is organized as follows. Section 2 introduces classical and quantum runtime, highlighting the implications for quantum program compilation. Section 3 examines high-level quantum programming and shows how multi-qubit gates naturally arise in such code. Section 4 then surveys the state-of-the-art decomposition algorithms used to break down these gates into executable instructions. Section 5 follows with a performance evaluation of these algorithms, focusing on CNOT count and circuit depth. Based on the benchmark data, Section 6 analyzes the trade-offs involved and introduces two practical quantum compilation profiles. Finally, Section 7 presents our conclusions and final remarks.
@@ 2. Classical and Quantum Runtime
Unlike classical code, which can be compiled once and reused with different inputs, a quantum circuit must be compiled separately for each specific input. This means that a quantum program designed to work with varying inputs can only be constructed at classical runtime–after all necessary parameters are known. As a result, the quantum portion of a classical-quantum program cannot be precompiled.
@@ 3. High-Level Quantum Programming
A controlled quantum gate call is the quantum analog of the classical if statement, where a gate is applied to the target qubits only if the control qubits are all in the state |$\ket{1}$. The key difference is that a controlled operation acts on the superposition and has the ability to create entanglement.
The code in Figure 2 illustrates how controlled operations can appear in high-level quantum programming. The function prepare takes a list of qubits alongside a list of measurement probabilities $[r_{k}]$ and a list of phase values $[\theta_{k}]$, and prepares the qubits in the state $\sum_{k}\sqrt{r_{k}}\,e^{i\theta_{k}}\ket{k}$. This function calls itself recursively, adding a control qubit at each recursion. Therefore, even though there is no direct call for controlled RY or P gates, the execution of this code creates several controlled operations.
{minted}
[frame=lines,fontsize=,breaklines]py def prepare( qubits: Quant, prob: ParamTree — list[float], amp: list[float] — None = None, ): if not isinstance(prob, ParamTree): prob = ParamTree(prob, amp) head, *tail = qubits RY(prob.value, head) if prob.is_leaf(): with around(X, head): P(prob.phase0, head) return P(prob.phase1, head) with around(X, head): ctrl(head, prepare)(tail, prob.left) ctrl(head, prepare)(tail, prob.right)
@@
Alongside controlled gate calls, Ket provides constructions that take advantage of the reversibility of quantum computation. A key example is the “with around” construction, which acts as a basis change–analogous to basis changes in linear algebra–for the inner scope. In the code of Figure 2, the with around construction is used to change the basis of the qubits from the one with control state |$\ket{1}$ to a basis where the control state is |$\ket{0}$. This construction also allows the compiler to remove some controlled gate calls, for the code from Figure 2, those associated with the X gate (Rosa et al., 2025).
@@ 4. Decomposition Algorithms
Pauli Gates V Chain C2X $n-2$ Clean $4n$ $6n$
@@
Single Aux Linear $1$ Clean/Dirty $8n$ $12n$
Log $1$ Clean $O(\log(n)^{3})$ $O(n\log(n)^{4})$
@@
Phase and Hadamard SU(2) Rewrite $1$ Clean $8n$ $12n$
Single Aux U(2) $1$ Clean $O(\log(n)^{3})$ $O(n\log(n)^{4})$
@@
In Ket, multi-qubit gates are constructed such that only multi-controlled versions of the gates X, Y, Z, RX, RY, RZ, P, and H arise from high-level programming. This constraint allows the quantum compiler to implement a streamlined and efficient set of decomposition algorithms. We categorize these gates into three distinct groups based on their decomposition requirements: (i) Pauli Gates (X, Y, and Z), a decomposition algorithm for one of these gates can be applied to the others with a simple basis change on the target qubit; (ii) Rotation Gates (RX, RY, and RZ), these belong to the special unitary group SU(2); (iii) Phase and Hadamard Gates (P and H), these belong to the broader unitary group U(2) [^4] All single-qubit gates belong to U(2), but some, like the rotation gates, are also in SU(2), a subgroup of U(2)..
Table 1 summarizes state-of-the-art quantum gate decomposition algorithms. Some of these algorithms require auxiliary qubits, which are additional qubits beyond those directly involved in the controlled gate. Depending on the algorithm, auxiliary qubits must either be initialized in the |$\ket{0}$ state (Clean auxiliary) or can be in any state (Dirty auxiliary). Regardless of their initial state, auxiliary qubits must be restored to their original state after decomposition so that they can be reused elsewhere and do not create new entanglements. Typically, decompositions using clean qubits are more efficient. The Ket compiler automatically manages the allocation and usage of auxiliary qubits when available (Rosa et al., 2024).
@@ 4.1. Pauli Gates
The Pauli gates are among the most commonly used quantum gates in algorithm design. A key insight for gate decomposition is that the decomposition of one Pauli gate can be adapted for the others. For example, in Figure 4(a), the Y and Z gates are defined using the X gate with a basis change. Consequently, the circuits for their controlled versions, shown in Figures 4(b) and 4(c), rely on a controlled X gate. This example explicitly illustrates the behavior of Ket’s compiler, which automatically applies this principle: only the decomposition of the Pauli X gate is explicitly defined, while the other Pauli gates are implemented through basis changes. In this paper, we evaluate only the performance of the X gate decomposition, as the decomposition of the other Pauli gates requires only an additional two or four single-qubit gates.
{minted}
[frame=lines,fontsize=,breaklines]py def myY(q): HS = cat(H, S) with around(HS, q): X(q) def myZ(q): with around(H, q): X(q) def myCY(c, t): ctrl(c, myY)(t) def myCZ(c, t): ctrl(c, myZ)(t)
@@
The Network decomposition (Nielsen and Chuang, 2010, p. 183) is one of the most efficient algorithms and can be applied to any multi-controlled gate. Figure 3 illustrates two variants of the algorithm for a 8-controlled Pauli X gate. Note that, in addition to the target and control qubits, this decomposition requires auxiliary qubits, all initialized to the |$\ket{0}$ state. For Pauli gate decomposition, we can reduces the number of auxiliary qubits by one or two compared to other gates.
@@ V Chain
The Linear algorithm relies on the application of a multi-controlled $2\pi$ rotation gate into an auxiliary qubit, as illustrated in Figure 6(a). This results in a multi-controlled Z gate with target and control qubits, which can be transformed into an multi-controlled X via a basis change at the target. The decomposition of the multi-controlled rotation is performed without auxiliary with the SU(2) Linear algorithm (Zindorf and Bose, 2024).
@@
The Linear Depth decomposition proposed by Da Silva and Park (2022), as the name suggests, results in a circuit with linear depth. This is an efficient algorithm that requires no auxiliary qubits and can be used for any quantum gate. We consider this algorithm as the fallback decomposition since it has no requirements to be applied to any U(2) gate. However, this algorithm results in a circuit with a quadratic number of CNOTs.
@@ 5. Algorithms Benchmark
The Compilation Time profile, as the name suggests, is optimized to reduce the classical execution time of the decomposition algorithms, which may also result in improved execution time for simulated quantum executions. This profile is also suitable for NISQ computers (Preskill, 2018) with quantum programs with up to 400 qubits. The Execution Time profile targets large-scale Fault-Tolerant quantum computers (Devitt et al., 2013; Google Quantum AI and Collaborators et al., 2025) with the potential to compute with millions of qubits.
@@
Figure 9 presents the number of CNOTs and circuit depth for multi-controlled Pauli gates. The data was obtained with the instruction “ctrl(c, X)(t)”, but for the other Pauli gates, there is only a constant number of single-qubit gates’ difference, which does not affect the presented data.
@@
Figure 10 presents the number of CNOTs and circuit depth for multi-controlled Rotation gates. The data was obtained using the instruction “ctrl(c, RX(pi))(t)44footnotetext: We set pi = 1.570796327., but for the other Rotation gates, there is only a constant number of single-qubit gates’ difference, which does not affect the presented data.
@@
Figure 11 presents the number of CNOTs and circuit depth for multi-controlled Phase and Hadamard gates. The data was obtained with the instruction “ctrl(c, H)(t)”, but for the Phase gates, there is only a constant number of CNOT gates’ difference, which does not significantly affect the data.
@@ 6. Results Analysis
Pauli Gates $1^{\text{th}}$ Network C2X $1^{\text{th}}$ Network C2X
@@
Rotation Gates $1^{\text{th}}$ Network C3X $1^{\text{th}}$ Network C2X
@@
Phase and Hadamard $1^{\text{th}}$ Network C3X $1^{\text{th}}$ Network C2X
@@ 7. Final Remarks
Another direction for future work is to analyze the performance of the decomposition algorithms in the context of non-Clifford gates (Aaronson and Gottesman, 2004), as these gates have a significant impact on quantum execution time when the program is encoded in a Quantum Error Correction code (Piveteau et al., 2021). Additionally, examining the impact of circuit optimization techniques, such as ZX-calculus (Kissinger and Van De Wetering, 2020), on the final circuit would be an interesting direction for further research.
Acknowledgements.
EID acknowledges the Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq through grant number 409673/2022-6 and ECRR acknowledges the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - CAPES, Finance Code 001.

### v2

@@ Abstract.
Quantum Computing, Quantum Programming, Quantum Compiler, Gate Decomposition, Ket
copyright: noneconference: XXIX Brazilian Symposium on Programming Languages; September 22–26, 2025; Recife, PE
@@ 1. Introduction
For programming languages and platforms that treat the quantum bit (qubit) as a first-class citizen in quantum programming, e.g., Q# (Svore et al., 2018) and Ket (Da Rosa and De Santiago, 2022), coding a quantum application involves invoking functions known as quantum gates. Quantum gates have no side effects on the classical state of the program; they only affect the quantum state and can induce superposition and entanglement (Nielsen and Chuang, 2010). Another class of functions, known as measurements, is used to extract information from the quantum state, returning it as classical data while causing the collapse of the quantum state, i.e., destroying the superposition.
Figure 1(a) illustrates a simple quantum program written in Python using the Ket quantum programming platform (Da Rosa and De Santiago, 2022). This code implements the Grover diffusion operator, a key component of Grover’s quantum search algorithm (Grover, 1996), utilizing the Hadamard (H), Pauli X, and Pauli Z gates. The term quantum gate, or more precisely quantum logical gate, is inspired by classical circuit and logic gates. Each quantum code has an equivalent quantum circuit; for example, the circuit in Figure 1(b) is equivalent to the code in Figure 1(a) applied for 4-qubits.
@@
def diffusion(q : Quant):
with around(cat(H, X), q):
with control(q[:-1]):
Z(q[-1])
@@
Figure 1. Grover diffusion operation implemented using Ket (LABEL:sub@fig:ket:diffusion), along with its corresponding high-level (LABEL:sub@fig:cirq:diffusion) and decomposed (LABEL:sub@fig:cirq:diffusion_decompose) quantum circuit for 4 qubits.
@@
The remainder of this paper is organized as follows. Section 2 introduces classical and quantum runtime, highlighting the implications for quantum program compilation. Section 3 examines high-level quantum programming and shows how multi-qubit gates naturally arise in such code. Section 4 then surveys the state-of-the-art decomposition algorithms used to break down these gates into executable instructions. Section 5 follows with a performance evaluation of these algorithms, focusing on CNOT count and circuit depth. Based on the benchmark data, Section 6 analyzes the trade-offs involved and introduces two quantum compilation profiles. Finally, Section 7 presents our conclusions and final remarks.
@@ 2. Classical and Quantum Runtime
Unlike classical code, which can be compiled once and reused with different inputs, a quantum circuit must be compiled separately for each specific input. This means that a quantum program designed to work with varying inputs can only be constructed at classical runtime–after all necessary parameters are known. As a result, the quantum portion of a classical-quantum program cannot be compiled in advance.
@@ 3. High-Level Quantum Programming
[…]
def prepare(
qubits: Quant,
prob: ParamTree | list[float],
amp: list[float] | None = None,
if not isinstance(prob, ParamTree):
prob = ParamTree(prob, amp)
head, *tail = qubits
RY(prob.value, head)
if prob.is_leaf():
with around(X, head):
P(prob.phase0, head)
return P(prob.phase1, head)
with around(X, head):
ctrl(head, prepare)(tail, prob.left)
ctrl(head, prepare)(tail, prob.right)
@@
[…]
@@ 4. Decomposition Algorithms
Pauli Gates V Chain C2X n2n-2 Clean $4n$ $6n$
@@
Single Aux Linear 11 Clean/Dirty $8n$ $12n$
Log 11 Clean $O(\log(n)^{3})$ $O(n\log(n)^{4})$
@@
Phase and Hadamard SU(2) Rewrite 11 Clean $8n$ $12n$
Single Aux U(2) 11 Clean $O(\log(n)^{3})$ $O(n\log(n)^{4})$
@@
In Ket, multi-qubit gates are constructed such that only multi-controlled versions of the gates X, Y, Z, RX, RY, RZ, P, and H arise from high-level programming. This constraint allows the quantum compiler to implement a streamlined and efficient set of decomposition algorithms. We categorize these gates into three distinct groups based on their decomposition requirements: (i) Pauli Gates (X, Y, and Z), a decomposition algorithm for one of these gates can be applied to the others with a simple basis change on the target qubit; (ii) Rotation Gates (RX, RY, and RZ), these belong to the special unitary group SU(2); (iii) Phase and Hadamard Gates (P and H), these belong to the broader unitary group U(2) [^4] All single-qubit gates are in U(2), with Rotation gates in SU(2) $\subset$ U(2)..
[…]
@@ 4.1. Pauli Gates
The Pauli gates are among the most commonly used quantum gates in algorithm design. A key insight for gate decomposition is that the decomposition of one Pauli gate can be adapted for the others. For example, in Figure 4(a), the Y and Z gates are defined using the X gate with a basis change. Consequently, the circuits for their controlled versions, shown in Figures 4(b) and 4(c), rely on a controlled X gate. This example explicitly illustrates the behavior of Ket’s compiler, which automatically applies this principle: only the decomposition of the Pauli X gate is explicitly defined, while the other Pauli gates are implemented through basis changes. In this paper, we evaluate only the performance of the X gate decomposition, as the decomposition of the other Pauli gates requires only an additional two single-qubit gates.
def myY(q):
with around(SD, q):
X(q)
def myZ(q):
with around(H, q):
X(q)
[…]
ctrl(c, myY)(t)
def myCZ(c, t):
ctrl(c, myZ)(t)
@@
[…]
@@ V Chain
The Linear algorithm relies on the application of a multi-controlled $2\pi$ rotation gate into an auxiliary qubit, as illustrated in Figure 6(a). This results in a multi-controlled Z gate with target and control qubits, which can be transformed into a multi-controlled X via a basis change at the target. The decomposition of the multi-controlled rotation is performed without auxiliary with the SU(2) Linear algorithm (Zindorf and Bose, 2024).
@@
The Linear Depth decomposition algorithm proposed by Da Silva and Park (2022), as the name suggests, results in a circuit with linear depth. This is an efficient algorithm that requires no auxiliary qubits and can be used for any quantum gate. We consider this algorithm as the fallback decomposition since it has no requirements to be applied to any U(2) gate. However, this algorithm results in a circuit with a quadratic number of CNOTs.
@@ 5. Algorithms Benchmark
The Compilation Time profile, as the name suggests, is optimized to reduce the classical execution time of the decomposition algorithms, which may also result in improved execution time for simulated quantum executions. The Execution Time profile targets large-scale Fault-Tolerant quantum computers (Devitt et al., 2013; Google Quantum AI and Collaborators et al., 2025) with the potential to compute with millions of qubits.
@@
Figure 9 presents the number of CNOTs and circuit depth for multi-controlled Pauli gates. The data was obtained with the instruction “ctrl(c, X)(t)”, but for the other Pauli gates, there is a difference of only a constant number of single-qubit gates, which does not affect the presented data.
@@
Figure 10 presents the number of CNOTs and circuit depth for multi-controlled Rotation gates. The data was obtained using the instruction “ctrl(c, RX(pi/2))(t)”, but for the other Rotation gates, there is a difference of only a constant number of single-qubit gates, which does not affect the presented data.
@@
Figure 11 presents the number of CNOTs and circuit depth for multi-controlled Phase and Hadamard gates. The data was obtained with the instruction “ctrl(c, H)(t)”, but for the Phase gates, there is a difference of only a constant number of single-qubit gates, which does not significantly affect the data.
@@ 6. Results Analysis
Pauli Gates $1^{\text{st}}$ Network C2X $1^{\text{st}}$ Network C2X
@@
Rotation Gates $1^{\text{st}}$ Network C3X $1^{\text{st}}$ Network C2X
@@
Phase and Hadamard $1^{\text{st}}$ Network C3X $1^{\text{st}}$ Network C2X
@@ 7. Final Remarks
Another direction for future work is to analyze the performance of the decomposition algorithms in the context of non-Clifford gates (Aaronson and Gottesman, 2004), as these gates have a significant impact on quantum execution time when the program is encoded in a Quantum Error Correction code (Piveteau et al., 2021). Additionally, examining the impact of circuit optimization techniques, such as ZX-calculus (Coecke and Duncan, 2011), on the final circuit would be an interesting direction for further research.
ARTIFACT AVAILABILITY
The authors declare that the research artifacts supporting the findings of this study are accessible at https://doi.org/10.5281/zenodo.16964846.
ACKNOWLEDGMENTS
ECRR acknowledges the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - CAPES, Finance Code 001; EID, JM, and ECRR acknowledges the Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq through grant number 409673/2022-6; JM, EID, and ECRR acknowledges the Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina - FAPESC through Project FAPESC TR nº 2024TR002672.